OpenStudy (anonymous):

solve: 5^(3x-4) = 15^(2x+1)

OpenStudy (anonymous):

Someone please explain how to do this. I know I have to use logs somehhow

OpenStudy (slaaibak):

I woud ln both sides: ln(5^(3x-4)) = ln(15^(2x+1)) then you get: (3x - 4) * ln(5) = (2x + 1) * ln(15) 3x - 4 = 2x* ln15/ln5 + ln15/ln5 x(3 - 2ln15/ln5) = 4 + ln 15/ln5 x = (4 + ln15/ln5)/(3 - ln 225 / ln 5)

OpenStudy (anonymous):

IS there not a way to get both sides of the equation to have the same base and then just set the exponents equal to each other?

OpenStudy (slaaibak):

hmmm, I suppose you could do this: 5^(3x-4) = 3^(2x + 1) * 5^ (2x + 1) Then, 3^(2x+1) = \[5^(\log_{5} 3)(2x+1)\] then it becomes: 5^(3x-4) = \[5^(2x +1 + (\log_{5} 3)(2x+1))\] then you can set it equal: 3x-4 = 2x +1 + (\log_{5} 3)(2x+1)

OpenStudy (anonymous):

Thanks :)